# Metastability and Ostwald Step Rule in the Crystallisation of Diamond and Graphite from Molten Carbon

### v1

@@
Comparing the qualitative features of the growing nuclei in direct simulations to FFS confirms that the chosen order parameters to define the flux interfaces do not bias the sampling towards unrealistic nucleation pathways. Figure 3 shows snapshots along the crystallisation pathway of graphite and diamond from unbiased simulations in which liquid carbon is cooled at 15 and 15.5 GPa. Diamond nucleates into compact crystallites with an isotropic aspect ratio, which justifies the spherical nucleus approximation often adopted in CNT. Similar behaviour was observed in the crystallisation of other tetrahedral systems, including silicon, germanium, and water ice.Li, Donadio, and Galli (2009); Li et al. (2011) This also means that the average estimate of $\gamma_{LS,D}$ is reliable. Conversely, graphite nucleation proceeds through the formation and stacking of small graphite patches which produce anisotropic crystallites elongated along the cross-plane axis of graphite (Figure 3(top). The growing graphite crystals are highly non-spherical and the free energy of the interface between the different facets and the liquid dictates their shape.
Employing the same order parameters used to analyse the crystallisation trajectories, we calculated the roughness of the interface between liquid carbon and the basal (0001) and primary prismatic (1000) planes of graphite in the two-phase simulations used to compute the melting line. From this roughness, it is possible to estimate the interfacial free energy $\gamma_{LS,G}$ for each surface, yielding 1.3±$\pm 0.08$ Jm-2 for the basal plane and $0.73\pm 0.07$ Jm-2 for the prismatic plane. The lower interfacial tension estimate of the primary prismatic facet is consistent with the findings of the direct MD and FFS simulations where graphite was found to nucleate through the formation of anisotropic crystallites elongated along the cross-plane axis of graphite that resemble the primary prismatic facet of graphite. The basal plane $\gamma_{LS,G}$ is in agreement with previous estimates from a continuum theory of carbon phases fitted to thermodynamic data (1.6 Jm-2).Umantsev and Akkerman (2010); Fried and Howard (2000) However, these theories cannot grasp the large differences among planes which dictate the nucleation pathways at the molecular scale.
Figure 4: The two contour plots show the logarithmic probability distribution of the size of clusters of carbon atoms selected according to their density and crystallinity order parameter in liquid carbon at 3650 K. In (a) the x-axis indicates the size of the biggest low-density cluster ($\rho<2.45$ g/cm3 within a sphere of 3 Å radius) at a pressure of 15 GPa, and in (b) of the biggest high-density cluster ($\rho>3.4$ g/cm3 within a sphere of 3 Å radius) at a pressure of 15.5 GPa. The dotted lines represent spontaneous nucleation pathways of graphite (a) and diamond (b). Panel (c) represents the one-step classical nucleation pathway of diamond at high pressure and the non-classical pathway that leads to the metastable crystallization of graphite at $P\lesssim 15$ GPa, which will eventually transform into diamond through a solid-solid phase transition.
@@
Previous considerations indicate that density is the main similarity parameter that drives the crystallisation of graphite or diamond from molten carbon since both transitions involve large density variations. At the tipping pressure of 15 GPa the crystallisation of graphite is accompanied by a density drop from 2.9 to 2.5 g/cm3, whereas when diamond is formed the density rises to 3.5 g/cm3. Such extreme density differences between the liquid and the solid phases affect nucleation pathways. Previous studies showed that, following Ostwald’s step rule, nucleation could be a two-step process in which density fluctuations with the formation of high-density amorphous or liquid aggregates precede crystalline ordering. This effect was observed during crystal nucleation both in supercooled single-component liquids and supersaturated solutions.Wolde and Frenkel (1997); ten Wolde and Frenkel (1999); Schilling et al. (2010); Hu and Tanaka (2022); Finney and Salvalaglio (2022) We observe a similar effect in the crystallisation of graphite. Figure 4(a) shows the probability that low-density clusters and crystalline nuclei of a given size form spontaneously in the liquid at 3650 K and 15 GPa. The shape of the distribution indicates that the crystallinity and low-density order parameters do not correlate. The occurrence of crystalline clusters without local low-density fluctuations can be explained as the system tends to form diamond-like small nuclei. However, spontaneous nucleation pathways to graphite, one of which is represented in the graph as a red dotted line, evolve through the formation of a low-density cluster that eventually develops into a crystalline nucleus. In the case of diamond nucleation instead, as shown in Figure 4(b), the density fluctuations producing high-density clusters are correlated to the crystallinity order parameter. Consequently, the crystalline nucleus is not preceded by a higher-density disorder precursor: local densification and ordering coincide. This difference further supports the evidence that graphite nucleation follows a lower-barrier free energy pathway, facilitated by local low-density fluctuations in liquid carbon even at the thermodynamic conditions at which diamond is the stable phase.
Conclusions
@@
Accurate and efficient machine learning potentials trained on density functional theory calculations, enabling a total of several microseconds of MD simulations of systems of several thousand atoms, provided us with an invaluable tool to explore the phase diagram of carbon near the GDL triple point and the intricate kinetics of homogenous nucleation of graphite and diamond from the melt. Whereas liquid carbon is often considered a glass former, we observed spontaneous crystallisation of both graphite and diamond in unbiased MD simulations in which the liquid is cooled rapidly at constant pressure. This suggested that liquid carbon at pressures between 5 and 30 GPa is not a glass former, and provided us with unbiased trajectories to examine the nucleation pathways at the molecular scale. We observed an unexpected wealth of physical phenomena related to homogeneous nucleation in a deceivingly simple monoatomic system such as liquid carbon. These phenomena are inherently connected to Ostwald’s step rule that leads to highly non-classical nucleation effects. In MD simulations, graphite crystallises spontaneously in the domain of thermodynamic stability of diamond up to 15 GPa due to the smaller density differences between liquid carbon and graphite and the lower liquid/solid interfacial free energy. In this range of pressures, graphite is a long-lived metastable state on the way to the crystallization of diamond.Bundy (1963b) Additionally, we found graphite nucleates through a non-classical two-step process where crystalline ordering is preceded by the formation of a low-density liquid region that further lowers the barrier to nucleation. Large differences in the interfacial free energy of different facets dictate the strongly anisotropic shape of growing graphite crystallites. Conversely, at pressures above 15 GPa, diamond nucleates through a classical one-step process, forming densely packed tetrahedral bonded crystalline spherical clusters. The two intrinsically different nucleation mechanisms of graphite and diamond are schematically represented in Figure 4c. Even if graphite nucleation is non-classical, a CNT fit of the nucleation rates yields physically meaningful parameters for both graphite and diamond, suggesting that even non-classical nucleation processes can be mapped on a CNT model.Zahn (2015)
@@
The insight into homogeneous nucleation from carbon melt provided by our simulations may help interpret inconsistencies among historical electrical and laser flash-heating experiments aimed at resolving the phase diagram of carbon near the GDL triple point. Depending on the details of these experiments and the recrystallization conditions, the system may remain trapped in metastable graphitic configurations. These observations may also impact the manufacturing of carbon-based materials such as synthetic diamonds and nanodiamonds at high pressure and high temperature.
Acknowledgements.
We are grateful to Gabriele C. Sosso for providing a critical assessment of the manuscript. This work was supported by the National Science Foundation under Grant No. 2053235.
Data Availability Statement
The data that support the findings of this study are openly available in [repository name] at http://doi.org/[doi], reference number [reference number].
@@ Methods
Molecular dynamics simulations were performed using the GPUMD v3.4 code with a NEP3 neuroevolution potentialFan et al. (2021, 2022) fitted to density functional theory (DFT) calculations in the local density approximation (LDA) for the exchange and correlation functional. The database of configurations, energies, forces, and virials was formerly generated to study amorphous carbon.Deringer and Csányi (2017) We conducted extensive tests to verify that the resulting NEP potential is suitable to model diamond and graphite. Besides the phase diagram shown in Figure 1, we verified that the potential reproduces the structural, vibrational, and elastic properties of diamond and graphite accurately (Table S1)These results suggest that, whereas LDA is the lowest-level approximation for exchange and correlation in DFT, it accurately describes the mechanical and thermodynamic properties of graphite, particularly the liquid-graphite phase coexistence. In the pressure range of interest, i.e. above 10 GPa, our predictions for the melting line of diamond are within 100 K of DFT calculations with GGA functionals.Wang, Scandolo, and Car (2005); Correa, Bonev, and Galli (2006)
Additionally, we verified the sensitivity of our predicted phase diagram and spontaneous crystallisation conditions upon the hyper-parameters and the training process of the NEP potential. We generated three more carbon NEP potentials varying several hyperparameters, including cutoff (cutoff radii for the radial and angular descriptor components), n_max (number of radial functions for the radial and angular descriptor components), basis_size (number of radial basis functions for the radial and angular descriptor components), and neuron (number of neurons in the hidden layer). Extended data Figure 1 shows that the melting line of graphite may shift by at most 100 K, whereas the melting line of diamond is insensitive to the different NEP parameterisations. Accordingly, crystallisation temperatures undergo variations up to 150 K, including the uncertainty inherent to the stochastic nature of this process. The highest pressure at which metastable graphite crystallises spontaneously varies between 13 and 15 GPa.
@@
MD simulations probing the properties of liquid carbon and spontaneous crystallisation were performed using the GPUMD package for cubic systems containing 4096 carbon atoms.Fan et al. (2022) The equations of motion were integrated with a timestep of 0.5 fs. The isobaric canonical ensemble (constant number of particles, pressure, and temperature - NPT) was controlled using the stochastic rescaling scheme with coupling times of 1 ps for the temperature and 5 ps for the pressure.Bussi, Donadio, and Parrinello (2007); Bernetti and Bussi (2020) Spontaneous crystallisation of diamond and graphite was observed in direct MD simulations of liquid carbon at constant pressure, in which the temperature was ramped from 5,000 K to 3,500 K in 25 ns, corresponding to a cooling rate of 60 K/ns.
We used phase coexistence simulations to determine the phase coexistence temperatures of liquid carbon with diamond and graphite.Morris et al. (1994); Correa, Bonev, and Galli (2006) These runs were set up by joining a diamond/graphite simulation cell with a liquid simulation cell equilibrated at 5000 K. Liquid was put in contact with the (100) facet of diamond and the (1000) facet of graphite. These simulations comprise systems of about 10,000 carbon atoms, thus avoiding issues with size effects. For each pressure, we ran a series of NPT simulations at temperatures chosen every $\Delta T=50~{}K$. We define the melting point as the temperature between two simulations at $T$ and $T+\Delta T$ in which the crystalline phase grows ($T$) and shrinks ($T+\Delta T$).
@@
.3 Nucleation rates from forward flux sampling including order parameters
@@
Nucleation rates of diamond and graphite are computed using our implementation of forward flux sampling (FFS) in LAMMPS, based on carbon NEP. In FFS, nucleation rate $R$ is given as the product of $\Phi_{\lambda_{0}}$, which is the initial flux rate of nucleation trajectory crossing the initial milestone $\lambda_{0}$, and $P(\lambda_{F}|\lambda_{0})$, which measures the probability for a trajectory starting from $\lambda_{0}$ and successfully reaching the final milestone $\lambda_{F}$. $\Phi_{\lambda_{0}}$ is obtained by $N_{0}/t_{0}V$, where $N_{0}$ (120) is the number of successful crossing collected at $\lambda_{0}$, $V$ is the volume of simulation cell, $t_{0}$ is the total time for $N_{0}$ crossings. $P(\lambda_{F}|\lambda_{0})=\prod_{i=1}^{n}P(\lambda_{i}|\lambda_{i-1})$, where $P(\lambda_{i}|\lambda_{i-1})=N_{i}/M_{i-1}$ is computed through conducting $M_{i-1}$ MD trial runs starting from the milestone $\lambda_{i-1}$ and collecting $N_{i}$ (120) successful crossing at the next adjacent interface $\lambda_{i}$. For each thermodynamic condition $T/P$, three independent FFS runs are carried out to obtain the geometric mean rate $\langle R\rangle=(\prod_{i=1}^{3}R_{i})^{1/3}$ and the standard error of $\ln\langle R\rangle$ is given by $\sqrt{\sum_{i=1}^{3}(\ln R_{i}-\ln\langle R\rangle)^{2}}/3$ (cite JCP 158 124501 (2023)).
@@
In FFS, the milestones $\lambda_{i}$ are defined based on the size of the largest crystalline cluster in the melt. Three types of carbon atoms, namely, liquid-like, diamond-like, and graphite-like, are differentiated through a local bond-order parameter $q_{n}$. A diamond-like carbon is defined as the carbon atom that has a $q_{6}>0.5$ (see Extended Data Figure 2a) and exactly 4 nearest neighbours, whereas a graphite-like carbon is the one that has a $q_{3}<-0.85$ (see Figure 2b) and exactly 3 nearest neighbours. These choices are made based on both the distributions of $q_{n}$ for the three phases and the characteristic local coordination in diamond and graphite. For the $q_{n}$ analysis, the cut-off distance in identifying the nearest neighbours in both diamond and graphite is set to be 1.8 Å. In determining the size of a crystalline cluster through the connectivity analysis, the same cut-off distance 1.8 Å is used for diamond, whereas a greater cut-off distance 3.5 Å is used for graphite, to account for the inter-layer spacing between neighbouring graphene sheets.
.4 Interfacial tension calculations
@@
$\mathrm{\gamma}\equiv\frac{k_{B}T}{2\pi\langle\sigma^{2}\rangle}\ln{\frac{L}{\xi}}$ (3)
where $T$ is the absolute temperature, $k_{B}$ is the Boltzmann constant, $\sigma^{2}$ is the mean squared fluctuation of molecules at the interface, $L$ is determined by the size along the $x$- or $y$- dimension (assuming $z$ is normal to the surface) and $\xi$ is the bulk correlation length.Cordeiro (2003) Using this relationship we are able to estimate the interfacial free energy differences between the liquid carbon and basal facet of graphite interface and liquid carbon and prismatic facet of graphite interface, using the phase coexistence MD simulations previously described. We set $L=38.046$ according to our simulation cell size, and $\xi=1.9$ corresponding to the first minimum of the radial distribution function of liquid carbon. $\sigma^{2}$ is calculated as: $\langle\sigma^{2}\rangle\equiv\langle(z-\langle z\rangle)^{2}\rangle$ where we selected atoms from the largest liquid carbon cluster belonging to one of the two interfaces in the periodic cell. The calculation is halted when the chosen interface becomes locally vertical (i.e., the approximate tangent plane contains $\hat{z}$) or makes contact with the other interface. We carried out the comparison between the two facets for three simulations each held at 10 GPa with varying temperatures (4600 K, 4700 K, and 4800 K).

### v2

@@ Results
Comparing the qualitative features of the growing nuclei in direct simulations to FFS confirms that the chosen order parameters to define the flux interfaces do not bias the sampling towards unrealistic nucleation pathways. Figure 3 shows snapshots along the crystallisation pathway of graphite and diamond from unbiased simulations in which liquid carbon is cooled at 15 and 15.5 GPa. Diamond nucleates into compact crystallites with an isotropic aspect ratio, which justifies the spherical nucleus approximation often adopted in CNT. Similar behaviour was observed in the crystallisation of other tetrahedral systems, including silicon, germanium, and water ice.Li, Donadio, and Galli (2009); Li et al. (2011) This also means that the average estimate of $\gamma_{LS,D}$ is reliable. Conversely, graphite nucleation proceeds through the formation and stacking of small graphite patches which produce anisotropic crystallites elongated along the cross-plane axis of graphite (Figure 3(top). The growing graphite crystals are highly non-spherical, and the free energy of the interface between the different facets and the liquid dictates their shape.
[…]
@@
Previous considerations indicate that density is the main similarity parameter that drives the crystallisation of graphite or diamond from molten carbon, since both transitions involve large density variations. At the tipping pressure of 15 GPa, the crystallisation of graphite is accompanied by a density drop from 2.9 to 2.5 g/cm3, whereas when diamond is formed, the density rises to 3.5 g/cm3. Such extreme density differences between the liquid and the solid phases affect nucleation pathways. Previous studies showed that, following Ostwald’s step rule, nucleation could be a two-step process in which density fluctuations with the formation of high-density amorphous or liquid aggregates precede crystalline ordering. This effect was observed during crystal nucleation both in supercooled single-component liquids and supersaturated solutions.ten Wolde and Frenkel (1997, 1999); Schilling et al. (2010); Hu and Tanaka (2022); Finney and Salvalaglio (2022) We observe a similar effect in the crystallisation of graphite. Figure 4(a) shows the probability that low-density clusters and crystalline nuclei of a given size form spontaneously in the liquid at 3650 K and 15 GPa. The shape of the distribution indicates that the crystallinity and low-density order parameters do not correlate. The occurrence of crystalline clusters without local low-density fluctuations can be explained by the fact that the system tends to form diamond-like small nuclei. However, spontaneous nucleation pathways to graphite, one of which is represented in the graph as a red dotted line, evolve through the formation of a low-density cluster that eventually develops into a crystalline nucleus. In the case of diamond nucleation, instead, as shown in Figure 4(b), the density fluctuations producing high-density clusters are correlated to the crystallinity order parameter. Consequently, the crystalline nucleus is not preceded by a higher-density disorder precursor: local densification and ordering coincide. This difference further supports the evidence that graphite nucleation follows a lower-barrier free energy pathway, facilitated by local low-density fluctuations in liquid carbon even at the thermodynamic conditions at which diamond is the stable phase.
Discussion
@@
Accurate and efficient machine learning potentials trained on density functional theory calculations, enabling a total of several microseconds of MD simulations of systems of several thousand atoms, provided us with an invaluable tool to explore the phase diagram of carbon near the GDL triple point and the intricate kinetics of homogeneous nucleation of graphite and diamond from the melt. Whereas liquid carbon is often considered a glass former, we observed spontaneous crystallisation of both graphite and diamond in unbiased MD simulations in which the liquid is cooled rapidly at constant pressure. This suggested that liquid carbon at pressures between 5 and 30 GPa is not a glass former, and provided us with unbiased trajectories to examine the nucleation pathways at the molecular scale. We observed an unexpected wealth of physical phenomena related to homogeneous nucleation in a deceivingly simple monoatomic system such as liquid carbon. These phenomena are inherently connected to Ostwald’s step rule, leading to highly non-classical nucleation effects. In MD simulations, graphite crystallises spontaneously in the domain of thermodynamic stability of diamond due to the smaller density differences between liquid carbon and graphite and the lower liquid/solid interfacial free energy. Depending on the details of the potential, metastable graphite may crystallize up to a pressure between 5 and 7.5 GPa above the graphite-diamond coexistence line. In this range of pressures, graphite is a long-lived metastable state on the way to the crystallisation of diamond.Bundy (1963b) Additionally, we found graphite nucleates through a non-classical two-step process where crystalline ordering is preceded by the formation of a low-density liquid region that further lowers the barrier to nucleation. Large differences in the interfacial free energy of different facets dictate the strongly anisotropic shape of growing graphite crystallites. Conversely, at higher pressure, diamond nucleates through a classical one-step process, forming densely packed tetrahedrally bonded crystalline spherical clusters. The two intrinsically different nucleation mechanisms of graphite and diamond are schematically represented in Figure 4c. Even if graphite nucleation is non-classical, a CNT fit of the nucleation rates yields physically meaningful parameters for both graphite and diamond, suggesting that even non-classical nucleation processes can be mapped on a CNT model.Zahn (2015)
@@
The insight into homogeneous nucleation from carbon melt provided by our simulations may help interpret inconsistencies among historical electrical and laser flash-heating experiments aimed at resolving the phase diagram of carbon near the GDL triple point. Depending on the details of these experiments and the recrystallisation conditions, the system may remain trapped in metastable graphitic configurations. These observations may also impact the manufacturing of carbon-based materials such as synthetic diamonds and nanodiamonds at high pressure and high temperature.
@@ Methods
Molecular dynamics simulations were performed using the GPUMD v3.6 code with a NEP3 neuroevolution potentialFan et al. (2021, 2022) fitted to density functional theory (DFT) calculations in the local density approximation (LDA) for the exchange and correlation functional (dubbed NEP@LDA) and a NEP4 potential developed based on the OptB88-vdW functional (NEP@OptB88-vdW).Klimeš, Bowler, and Michaelides (2010); Rowe et al. (2020) For both potentials, the databases of configurations, energies, forces, and virials were formerly generated to study diverse carbon systems, including graphite, diamond, low and high-density carbon, and liquid carbon.Deringer and Csányi (2017); Rowe et al. (2020) We conducted extensive tests to verify that the resulting NEP potentials are suitable for modelling diamond and graphite. Besides the phase diagrams shown in Figure 1 and S2, we verified that the NEP@LDA reproduces the structural, vibrational, and elastic properties of diamond and graphite accurately (Table S1). These results suggest that, whereas LDA is the lowest-level approximation for exchange and correlation in DFT, it accurately describes the mechanical and thermodynamic properties of graphite, particularly the liquid-graphite phase coexistence. In the pressure range of interest, i.e., above 10 GPa, our predictions for the melting line of diamond are within 100 K of DFT calculations with GGA functionals.Wang, Scandolo, and Car (2005); Correa, Bonev, and Galli (2006)
Additionally, we verified the sensitivity of our predicted phase diagram and spontaneous crystallisation conditions upon the hyperparameters and the training process of the NEP@LDA potentials. We generated three more carbon NEP@LDA potentials varying several hyperparameters, including cutoff (cutoff radii for the radial and angular descriptor components), n_max (number of radial functions for the radial and angular descriptor components), basis_size (number of radial basis functions for the radial and angular descriptor components), population size, and number of generation. Figure S1 shows that the melting line of graphite may shift by at most 100 K, whereas the melting line of diamond is insensitive to the different NEP parameterisations. Accordingly, crystallisation temperatures undergo variations up to 150 K, including the uncertainty inherent to the stochastic nature of this process. The highest pressure at which metastable graphite crystallises spontaneously varies between 13 and 15 GPa. The hyperparameters for training the NEP@LDA and NEP@OptB88 models are reported in Table S2.
@@
MD simulations probing the properties of liquid carbon and spontaneous crystallisation were performed using the GPUMD package for cubic systems containing 4096 carbon atoms.Fan et al. (2022) The equations of motion were integrated with a timestep of 0.5 fs. We have verified that with a 0.5 fs timestep, MD simulations of liquid carbon at $\sim 4500$ K conserve the energy within 1 meV/atom over 100 ps (Figure S6). The isobaric canonical ensemble (constant number of particles, pressure, and temperature - NPT) was controlled using the stochastic rescaling scheme with coupling times of 1 ps for the temperature and 5 ps for the pressure.Bussi, Donadio, and Parrinello (2007); Bernetti and Bussi (2020) Spontaneous crystallisation of diamond and graphite was observed in direct MD simulations of liquid carbon at constant pressure, in which the temperature was ramped from 5,000 K to 3,500 K in 25 ns, corresponding to a cooling rate of 60 K/ns. These simulations were initiated from liquid models equilibrated at 5000 K for at least 5 ns. The structural features of these liquids are characterised by the radial distribution functions and the coordination numbers reported in Figures S4 and S5.
.3 Calculation of the phase boundaries
[…]
To compute the graphite-diamond phase boundary and to refine the solid-liquid coexistence lines, we integrated the Gibbs-Duhem equationKofke (1993) by finite differences in $\beta$:
$\frac{dp}{d\beta}=-\frac{\Delta h}{\beta\Delta v}$ (3)
where $\beta=1/k_{B}T$, $\Delta h$ is the difference in molar enthalpy between the two phases, and $\delta v$ is the difference in molar volume. For the graphite-liquid coexistence line, we integrated Eq. 3 starting from the triple point with a fixed $\Delta\beta=0.05$ eV-1. At each step, $\Delta h$ and $\Delta v$ are estimated as ensemble averages over the second half of 100 ps NPT runs.
@@
.4 Nucleation rates from forward flux sampling and order parameters
@@
Nucleation rates of diamond and graphite are computed using our implementation of forward flux sampling (FFS) in LAMMPS, based on carbon NEP. In FFS, nucleation rate $R$ is given as the product of $\Phi_{\lambda_{0}}$, which is the initial flux rate of nucleation trajectory crossing the initial milestone $\lambda_{0}$, and $P(\lambda_{F}|\lambda_{0})$, which measures the probability for a trajectory starting from $\lambda_{0}$ and successfully reaching the final milestone $\lambda_{F}$. $\Phi_{\lambda_{0}}$ is obtained by $N_{0}/t_{0}V$, where $N_{0}$ (120) is the number of successful crossings collected at $\lambda_{0}$, $V$ is the volume of the simulation cell, $t_{0}$ is the total time for $N_{0}$ crossings. $P(\lambda_{F}|\lambda_{0})=\prod_{i=1}^{n}P(\lambda_{i}|\lambda_{i-1})$, where $P(\lambda_{i}|\lambda_{i-1})=N_{i}/M_{i-1}$ is computed through conducting $M_{i-1}$ MD trial runs starting from the milestone $\lambda_{i-1}$ and collecting $N_{i}$ (120) successful crossing at the next adjacent interface $\lambda_{i}$. For each thermodynamic condition $T/P$, three independent FFS runs are carried out to obtain the geometric mean rate $\langle R\rangle=(\prod_{i=1}^{3}R_{i})^{1/3}$ and the standard error of $\ln\langle R\rangle$ is given by $\sqrt{\sum_{i=1}^{3}(\ln R_{i}-\ln\langle R\rangle)^{2}}/3$.Zhao and Li (2023)
@@
In FFS, the milestones $\lambda_{i}$ are defined based on the size of the largest crystalline cluster in the melt. Three types of carbon atoms, namely, liquid-like, diamond-like, and graphite-like, are differentiated through a local bond-order parameter $q_{n}$.Li, Donadio, and Galli (2009) Similar to the Steinhardt bond order parameter,Steinhardt, Nelson, and Ronchetti (1983) $q_{n}$ is rotationally invariant; however, as a local order parameter, $q_{n}$ reflects the inherently local character of nucleation. A diamond-like carbon is defined as the carbon atom that has a $q_{6}>0.5$ (see Figure S8) and exactly 4 nearest neighbours, whereas a graphite-like carbon is the one that has a $q_{3}<-0.85$ (see Figure S8) and exactly 3 nearest neighbours. These choices are made based on both the distributions of $q_{n}$ for the three phases and the characteristic local coordination in diamond and graphite. For the $q_{n}$ analysis, the cut-off distance in identifying the nearest neighbours in both diamond and graphite is set to be 1.8 Å. In determining the size of a crystalline cluster through the connectivity analysis, the same cut-off distance 1.8 Å is used for diamond, whereas a greater cut-off distance 3.5 Å is used for graphite, to account for the inter-layer spacing between neighbouring graphene sheets.
.5 Interfacial tension calculations
@@
$\mathrm{\gamma}\equiv\frac{k_{B}T}{2\pi\langle\sigma^{2}\rangle}\ln{\frac{L}{\xi}},$ (4)
where $T$ is the absolute temperature, $k_{B}$ is the Boltzmann constant, $\sigma^{2}$ is the mean squared fluctuation of molecules at the interface, $L$ is determined by the size along the $x$- or $y$- dimension (assuming $z$ is normal to the surface) and $\xi$ is the bulk correlation length.Cordeiro (2003) Using this relationship, we are able to estimate the interfacial free energy differences between the liquid-graphite (basal facet) and the liquid-graphite (prismatic facet), using the phase coexistence MD simulations used to compute the liquid-solid phase boundary. We set $L=38.046$ according to our simulation cell size, and $\xi=1.9$ corresponding to the first minimum of the radial distribution function of liquid carbon. $\sigma^{2}$ is calculated as: $\langle\sigma^{2}\rangle\equiv\langle(z-\langle z\rangle)^{2}\rangle$ where we selected atoms from the largest liquid carbon cluster belonging to one of the two interfaces in the periodic cell. The calculation is halted when the chosen interface becomes locally vertical (i.e., the approximate tangent plane contains $\hat{z}$) or makes contact with the other interface. We carried out the comparison between the two facets for three simulations, each held at 10 GPa with varying temperatures (4600 K, 4700 K, and 4800 K).
Data Availability Statement
The data that support the findings of this study are available open-source at http://doi.org/10.5281/zenodo.15567723. The implementation of forward flux sampling in LAMMPS is available at https://github.com/WanyuZhao/FFS-LAMMPS. Source data are provided with this paper.
Acknowledgements.
We are grateful to Gabriele C. Sosso for providing a critical assessment of the manuscript. This work was supported by the National Science Foundation under Grants No. 2053235 (DD) and 2053330 (TL).
Author Contributions Statement
DD and TL conceived the study. DD and WZ conducted the molecular dynamics simulations. SC trained the machine learning models. WZ, MLB, SC, DD, and TL analysed the results. DD drafted the manuscript. All authors contributed to editing and finalizing the manuscript.
Competing Interests Statement
All authors declare no competing interests.